OpenStudy (anonymous):

The x intercept of a line is -5 and the y intercept of the line is -2. Whats the equation of this line?

OpenStudy (phoenixfire):

You have two points:(-5, 0) <- x intercept and (0, -2) <- y intercept

OpenStudy (phoenixfire):

Do you know how to find the slope of a line from two points?

OpenStudy (anonymous):

y = -5/2x-5 y=2/5x+2 y=5/2x-2 y=-2/5x-2

OpenStudy (anonymous):

im pretty sure its the third one because ive narrowed it down to those choices is c correct?

OpenStudy (anonymous):

It wouldn't be the third one because the slope has to be negative.

OpenStudy (anonymous):

ohh dang it! okay yes your right then its A right?

OpenStudy (anonymous):

It would be the last one. The slope is correct and the y intercept is correct.

OpenStudy (anonymous):

oh really? i thought for sure it was A could you explain it?

OpenStudy (phoenixfire):

Slope:\[m={{y_1-y_0} \over {x_1-x_0}}={{0-(-2)} \over {-5-0}}={2 \over -5}=-{2 \over 5}\] y-intercept is -2. so y=mx+c becomes: \[y=-{2 \over 5}x-2\]

OpenStudy (anonymous):

oh i feel dumb! I get it haha

OpenStudy (anonymous):

could you guys help me with one more question?

OpenStudy (anonymous):

What is the slope-intercept form of the equation y – 7 = –5/2 (x + 4)?

OpenStudy (anonymous):

Rearrange it for y, the slope would be -5/2

OpenStudy (anonymous):

ok so would it be y – 2 = –5/2(x + 2) or y = –5/2x – 3

OpenStudy (anonymous):

the second one.

OpenStudy (anonymous):

okay cool thanks!

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

what about this question What is the slope of the line passing through the points (2, 7) and (–1, 3)?

OpenStudy (anonymous):

4/3

OpenStudy (anonymous):

are you sure it isnt 3/4 tho?

OpenStudy (anonymous):

yep, its 4/3

OpenStudy (anonymous):

okay cool thanks alot you are a big help!

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

best response?